19 Athletic Sports, Inc., produces high-quality sports equipment. The companys Racket Division manufactures two tennis rackets -the Standard and the Deluxe-that are widely used in amateur play. Selected information on the rackets is given below Standard S47.70 Deluxe Selling price per racket Variable expenses per racket 84.00 Production Selling S21.40 S2.45S 28.10 5.50 Sales in units over the past two months have been as follows; Deluxe 3,000 3,500 Standard April May 5,000 4,000 All sales are made through the companys own retail outlets. The Racket Division has the following fixed costs Fixed production costs Advertising expense Administrative salaries Per Month $107,000 101,000 50,000 S 258,000 Total Compute the Racket Divisions break-even point in dollar sales for April S 460,185.43 S 430,000.00 $ 467,920.13 S 469,090.91 S 516,000.00
Answer #1

Answer:

Break Even Point (in Dollar Sales) = Fixed Cost / Contribution Margin Ratio

Standard Racket:
Contribution Margin per Unit = Selling Price per Unit – Variable Cost per unit
Standard’s Variable Cost per unit = $21.40 + $2.45 = $23.85
Standard’s Contribution Margin per Unit = $47.70 - $23.85
Standard’s Contribution Margin per Unit = $23.85

Weight of Standard Racket Sale for the month of April = 5,000 / 8,000 = 62.50%

Deluxe Racket:
Contribution Margin per Unit = Selling Price per Unit – Variable Cost per unit
Deluxe’s Variable Cost per unit = $28.10 + $5.50 = $33.60
Deluxe’s Contribution Margin per Unit = $84.00 - $33.60
Deluxe’s Contribution Margin per Unit = $50.40

Weight of Deluxe Racket Sale for the month of April = 3,000 / 8,000 = 37.50%

Weighted Contribution Margin for the month of April = (0.625 * 23.85) + (0.375 * 50.40)
Weighted Contribution Margin for the month of April = 14.90625 + 18.90
Weighted Contribution Margin for the month of April = $33.80625

Weighted Sales = (0.625 * 47.70) + (0.375 * 84.00)
Weighted Sales = 29.8125 + 31.50 = $61.3125

Weighted Contribution Margin Ratio = 33.80625 / 61.3125 * 100
Weighted Contribution Margin Ratio = 0.551376

Break Even Point (in Dollar Sales) = 258,000 / 0.551376
Break Even Point (in Dollar Sales) = $467,920.28